# A Million Three-body Binaries Caught by Gaia

### v2

@@ A.2 The SMA/Eccentricity Distribution
Prior to Y. B. Ginat & H. B. Perets (2024); D. Atallah et al. (2024), the joint SMA ($a$) and eccentricity ($\epsilon$) distribution of binaries formed exclusively through 3BBF were not known (S. J. Aarseth & D. C. Heggie, 1976b; J. Goodman & P. Hut, 1993; N. Ivanova et al., 2005). Y. B. Ginat & H. B. Perets (2024); D. Atallah et al. (2024) found that 3BBF features a constant, thermal distribution ($f_{F}(\epsilon)=2\epsilon$; see J. H. Jeans 1919) for binaries with an SMA $a<R_{1}$ and an increasingly superthermal distribution for all SMA in the regime $a>R_{1}$. There is a slight subthermal region that peaks between $10^{-0.75}$ and $1$. This subthermal region plays an outsized role in the eccentricity of tight binaries because the formation rate of binaries with $a<R_{1}$ peaks when $a\approx R_{1}$. Remarkably, the 2D $(\epsilon,a/R_{1})$ distribution exhibits (near-) scale-free behavior for regions of the parameter space corresponding to $\chi_{1}=R_{1}/(2b_{90})\gtrsim 10$. Thus, independent of the 3BBF probability, $P_{F}(\chi_{1}>10)$, the 2D probability density function $f(\epsilon,a/R_{1})\,d\epsilon\,da/R_{1}$ may be semi-analytically approximated with ease.
@@
$\frac{d^{4}\Gamma}{drdkd\epsilon da}=36c_{0}\sqrt{\frac{3G\langle m\rangle}{\pi b^{7}}}\left(\frac{M_{\rm cl}}{\langle m\rangle}\right)^{5/6}k_{1}^{2}k^{2}e^{-k^{3}}\frac{(r/b)^{2}}{(1+(r/b)^{2})^{61/12}}f_{F}\left(\epsilon,\frac{a}{k_{1}r_{\rm sep}(r)}\right).$
@@
[…]
and $x=a/R_{1}$ for integration convenience. In the final portion of the calculation (§ D), the substitution $r\rightarrow r_{t}\tilde{r}$ is applied to limit the integration boundary within the tidal radius of the Plummer cluster; $\tilde{r}$ is the radial displacement normalized by the tidal radius of the cluster. Equation (A10) is the “local creation rate” and we will not consider any other means to form binaries, whether through 3BBF outside of a Plummer cluster or other binary formation mechanisms. The “creation rate” is now written
@@ A.4 Cluster Evolution
$\tau_{\rm cl}(N_{\rm cl0},R_{G})=0.13N_{\rm cl0}^{3/4}\frac{R_{G}}{V_{G}},$
$r_{\rm h}(M_{\rm cl},R_{G})=0.66\zeta^{2/3}r_{t}(M_{\rm cl},R_{G})\,\frac{\ln\left[\Lambda(N_{\rm cl})\right]^{2/3}}{N_{\rm cl}^{1/6}}\left(1-\left(1-\tilde{\tau}\right)^{11/3}\right)^{2/3},$
@@
Figure 12 displays a host of cluster properties—born on circular orbits at $R_{G}=8\,\rm{kpc}$—as they evolve through their cluster lifetime, $\tilde{\tau}=\tau/\tau_{\rm cl}$. The cluster is dissolved when $\tilde{\tau}=1$. For example, a cluster born with $M_{\rm cl0}=10^{3.3}\,M_{\odot}$ will have a half-mass radius of $r_{h}\approx 1.5\,\rm{pc}$ after shedding half its initial mass through relaxation and Galactic tides. All the cluster properties are in qualitative agreement with the set of cluster $M_{\rm cl},r_{\rm h}$ observations compiled in M. R. Krumholz et al. (2019).
[…]
A.5 Escape & Disruption in a Stellar Cluster
Two criteria are required for a 3BB to escape a cluster: (i) the center-of-mass (CoM) velocity of the new binary must exceed the local escape speed of the cluster and (ii) the binary must not undergo any strong (ionizing) encounters en route out of the cluster’s gravity well. By definition, wide binaries are very easy to disrupt and so this calculation should be done precisely as to not overestimate their survival rate.
Two key equations will be necessary for the calcuating the survival and escape probabilities along the approximately radial path the binary takes, the first being the standard Maxwellian velocity distribution
$f_{M}(v,\sigma_{b})=\sqrt{\frac{2}{\pi}}\frac{v^{2}}{\sigma_{b}^{3}}e^{-v^{2}/(2\sigma_{b}^{2})},\ \sigma_{b}^{2}(\tilde{r})=\sigma^{2}(\tilde{r})+\frac{Gm}{18a}$ (A14)
where $\sigma_{b}$ is the velocity dispersion of the new binary at $\tilde{r}$. Assuming 3BBF between equal masses, the mean-square velocity kick imparted to the new binary is $Gm/(6a)$. The second equation is the equal-mass ionization cross section (P. Hut & J. N. Bahcall, 1983)
$\Sigma_{\rm ion}=\frac{20\pi}{3}\frac{Gm}{v_{\rm rel}^{2}}a$ (A15)
where $v_{\rm rel}=\sqrt{v_{\rm b}^{2}+v^{2}}=\sqrt{12\sigma^{2}(\tilde{r})+v^{2}}$ is the relative velocity between the target binary and a single star. We only consider the fraction of 3BB’s formed with a CoM velocity greater than the local escape speed and make the simplifying assumption that their velocity is always equal to the local escape speed and on a radial trajectory. Larger velocities act to suppress the ionization rate, but the fraction of 3BBF’s with CoM velocities greater than $1.2v_{\rm esc}$ is vanishingly small, and so we choose to ignore this subtly in favor of accounting for the slowing of the binary as it exits the cluster.
Employing $\Sigma$ for a trivial “N-$\Sigma$-V” calculation, the ionization rate is
$\Gamma_{\rm ion}$ $=\left\langle n\,\Sigma\,v_{\rm rel}\right\rangle=\frac{20\pi}{3}nGma\left\langle\frac{1}{v_{\rm rel}}\right\rangle,$ (A16)
[…]
and the expected number of ionizing encounters for an escaping binary becomes
$\lambda_{\rm ion,out}(r,M_{\rm cl},b,a)=\int^{r_{t}}_{r}\Gamma_{\rm ion}(r)\frac{dr}{v_{b}}=\frac{5\sqrt{3}\,\alpha_{0}\,}{2}\frac{a}{b}\left[\frac{\pi}{2}-f_{\lambda}(r)\right],$ (A17)
$f_{\lambda}(r)=\frac{r/b}{1+(r/b)^{2}}+\tan^{-1}(r/b)$
where $r$ is the birth location of the 3BB and where it begins its getaway. The fraction of 3BB’s paired with CoM velocity $v_{b}>v_{\rm esc}(\tilde{r})$ is
$P_{\rm esc}(\tilde{r},M_{\rm cl},b,a)=2\sqrt{\frac{6}{\pi}}\frac{\sigma}{\sigma_{b}}e^{-6(\sigma/\sigma_{b})^{2}}+\rm{Erfc}\left(\sqrt{6}\frac{\sigma}{\sigma_{b}}\right),\ \ \sigma_{b}^{2}=\sigma^{2}+\frac{Gm}{18a}.$ (A18)
Assuming a Poison process, the probability of a 3BB securing a clean escape with absolutely no ionizing encounters is
$P_{\rm ion,out}(\tilde{r},M_{\rm cl},b,a)=e^{-\lambda_{\rm ion,out}}$ (A19)
There are two escape trajectories to consider: (i) the binary travels on a radial path out of the cluster, Equation (A19), and (ii) the binary travels on a radial path inwards through the core of the cluster before continuing out of the cluster. Binaries taking an “inward” path have a survival probability equivalent to $P_{\rm ie,in}\approx P_{\rm ie,out}P_{\rm 0\rightarrow r}$. Here, $P_{\rm 0\rightarrow r}$ is the probability of no binary ionization after passing twice through the region spanning $0\rightarrow r$. Stated mathematically,
$P_{\rm ion,in}=P_{\rm ion,out}e^{-\lambda_{\rm ion,in}},\ \ \lambda_{\rm ion,in}=5\sqrt{3}\,\alpha_{0}\frac{a}{b}f_{\lambda}(r)$ (A20)
with the total probability of a clean escape
$P_{\rm ie}\left(\tilde{r},M_{\rm cl},b,a\right)$ $=P_{\rm esc}P_{\rm ion},\ \ P_{\rm ion}=\frac{e^{-\lambda_{\rm ion,out}}}{2}\left(1+e^{-\lambda_{\rm ion,in}}\right),$ (A21)
where we take the average of the two possibilities.
@@ Appendix B Star Formation History and Radial Diffusion in the Milky Way’s Disks
Up until § A.5, we only considered dynamics internal to toy Milky Way stellar clusters. To estimate the number and orbital element distributions of 3BB’s in the solar neighborhood, it is critical to model the relationship between the clusters hosting 3BBF, the MW disk, the cluster birth time, $t$, and Galactocentric radius, $R_{G}$, in the Milky Way thin and thick disks, and the journey (i.e., angular momentum diffusion) the few 3BB’s managing to escape their home clusters must take to reach a comparatively small spatial region within $\sim 200\,\rm{pc}$ of the Sun.
@@ B.1 Milky Way Disk Evolution
$SFH_{\rm ha}(t)=\begin{cases}\frac{e^{-t/\tau_{SFR}}}{\tau_{SFR}\,\left(1-e^{-\tau_{ha}/\tau_{SFR}}\right)},&t<\tau_{ha}\\ 0,&t>\tau_{ha}\end{cases}$
@@
where $R_{d0}=2.3\,\rm kpc$ is the disk scale-length, $R_{G}$ is the Galactocentric distance of the cluster’s orbit and the starting point from which escaping 3BBs begin their disk migration, $t$ is the time separation between the MW’s birth (assumed to occur $\tau_{\rm MW}=12\,\rm Gyrs$ ago) and the cluster’s birth, $\tau$ is the time between cluster birth and formation of the binary, $t^{\prime}=\tau_{\rm MW}-\tau-t$ is the lookback time since a 3BB formed in a cluster, and $R_{G}^{\prime}$ is the location of the binary at $t^{\prime}$. Additionally, $\tau_{SFR}=6.8\,\rm Gyrs$ is the star formation timescale of the thick disk, $\tau_{ha}=6\,\rm Gyrs$ is the time when new cluster formation is “turned off” in the thick disk, $\tau_{m}=6\,\rm Gyrs$ is a timescale controlling the diffusion strength in the disk, and $c_{1}=2.48\,\rm{kpc}$ is the diffusion scale length. For simplicity, the latter two quantities and the form of the diffusion distribution, $D_{ha}$, are assumed identical to that of the thin disk models given by N. Frankel et al. (2020).
@@
$SFH_{\rm la}(t,R_{G})=$$\begin{cases}\frac{y\,e^{y(\tau_{\rm MW}-\tau_{la}-t)/\tau_{SFR}}}{\tau_{\rm SFR}\left(1-e^{-y\,\tau_{la}/\tau_{SFR}}\right)},&\tau_{ha}<t<\tau_{MW}\\ 0,&\rm{otherwise}\end{cases}$ (B2)
@@
where $\tau_{la}=6\,\rm\,Gyr$ and $\tau_{SFR}=1\,\rm\,Gyr$. T. Wagg et al. (2022) constructed their Milky Way model assuming a total stellar mass $M_{d}=2.6\times 10^{10}M_{\odot}$ in each of the thick and thin disks and a bulge mass $M_{b}=0.9\times 10^{10}M_{\odot}$ (T. C. Licquia & J. A. Newman, 2015). To incorporate the stellar mass contribution from the bulge, we treat the bulge of the Milky Way as being built up from the combined stellar diffusion in the thick and thin disks, applying all the mass of the bulge to the older thick disk. The number of stars in each present-day disk are $N_{\rm ha}=\frac{M_{d}+M_{b}}{m}=7\times 10^{10}$ and $N_{\rm la}=\frac{M_{d}}{m}=5.2\times 10^{10}$, respectively. Combining the two disk distributions, our evolving MW disk model is

### v3

@@ A.2 The SMA/Eccentricity Distribution
Prior to Y. B. Ginat & H. B. Perets (2024); D. Atallah et al. (2024), the joint SMA ($a$) and eccentricity ($\epsilon$) distribution of binaries formed exclusively through 3BBF were not known (S. J. Aarseth & D. C. Heggie, 1976; J. Goodman & P. Hut, 1993; N. Ivanova et al., 2005). Y. B. Ginat & H. B. Perets (2024); D. Atallah et al. (2024) found that 3BBF features a constant, thermal distribution ($f_{F}(\epsilon)=2\epsilon$; see J. H. Jeans 1919) for binaries with an SMA $a<R_{1}$ and an increasingly superthermal distribution for all SMA in the regime $a>R_{1}$. There is a slight subthermal region that peaks between $10^{-0.75}$ and $1$. This subthermal region plays an outsized role in the eccentricity of tight binaries because the formation rate of binaries with $a<R_{1}$ peaks when $a\approx R_{1}$. Remarkably, the 2D $(\epsilon,a/R_{1})$ distribution exhibits (near-) scale-free behavior for regions of the parameter space corresponding to $\chi_{1}=R_{1}/(2b_{90})\gtrsim 10$. Thus, independent of the 3BBF probability, $P_{F}(\chi_{1}>10)$, the 2D probability density function $f(\epsilon,a/R_{1})\,d\epsilon\,da/R_{1}$ may be semi-analytically approximated with ease.
@@
[…]
@@
and $x=a/R_{1}$ for integration convenience. In the final portion of the calculation (§ D), the substitution $r\rightarrow r_{t}\tilde{r}$ is applied to limit the integration boundary within the tidal radius of the Plummer cluster; $\tilde{r}$ is the radial displacement normalized by the tidal radius of the cluster. Equation (A10) is the “local creation rate” and we will not consider any other means to form binaries. The “creation rate” is now written
@@ A.4 Cluster Evolution
[…]
@@
Figure 12 displays a host of cluster properties—born on circular orbits at $R_{G}=8\,\rm{kpc}$—as they evolve through their cluster lifetime, $\tilde{\tau}=\tau/\tau_{\rm cl}$ and fully dissolved when $\tilde{\tau}=1$. For example, a cluster born with $M_{\rm cl0}=10^{3.3}\,M_{\odot}$ will have a half-mass radius of $r_{h}\approx 1.5\,\rm{pc}$ after shedding half its initial mass through relaxation and Galactic tides. All the cluster properties are in qualitative agreement with the set of cluster $M_{\rm cl},r_{\rm h}$ observations compiled in M. R. Krumholz et al. (2019).
Figure 12: Assumed semi-analytic relations between cluster properties—initial mass $M_{\rm cl0}$, half-mass radius $r_{h}$, and density $\rho_{h}$ within the half-mass radius, and age $\tilde{\tau}$ normalized by cluster lifespan—according to our compilation of the UEE from M. Gieles et al. (2011)—Equations (A12). In the left column, the solid curves show the $r_{h}$$M_{\rm cl0}$ relation (top) and $\rho_{h}$$M_{\rm cl0}$ relation (bottom) at several cluster ages indicated by color. The time evolution of $r_{h}$, $\rho_{h}$ and $M_{\rm cl}$ for several specific choices of initial cluster mass $M_{\rm cl0}$ are indicated by the dashed curves in both the left and right columns. The early truncation of the dashed yellow and light blue curves prior to $\tilde{\tau}=1$ (cluster dissolution) reflect our choice to halt cluster evolution once there are fewer than 10 bodies, below which these semi-analytic approximations break down. All clusters shown are assumed to evolve on a circular orbit about the Milky Way center at an equivalent Galactocentric radius to the sun ($R_{G}=8\,\rm{kpc}$).
A.5 Survival & Disruption in a Stellar Cluster
Two criteria are required for a 3BB to escape a cluster: (i) the center-of-mass (CoM) velocity of the new binary must exceed the local escape speed of the cluster and (ii) the binary must not undergo any strong (ionizing) encounters en route out of the cluster’s gravity well. Unfortunately, the fraction of three-body encounters in which three unbound bodies encounter each other within a small volume, whilst simultaneously occupying the “tail” of a cluster’s velocity distribution, is vanishingly small. In fact, the outgoing binary CoM velocities from the 3UB simulations generated in § A.1 are approximately described by a Maxwellian with a velocity dispersion ${\sim}70\%$ (smaller) than the median of the initial velocity dispersion. It is thus safe to assume that newly formed 3BBs do not immediately escape upon assembly.
Instead of the “immediate escape” channel, we will consider the fraction of 3BBs which form in a cluster and survive within it until the time of the cluster’s absolute dissolution. Our calculation employs the geometric cross-section, $\Sigma=\pi a^{2}$, rather than the ionization cross-section (P. Hut & J. N. Bahcall, 1983; D. C. Heggie & P. Hut, 1993) as we are only interested in binaries which do not undergo a single encounter over the rest of the cluster lifetime. We emphasize that this cross-section is strictly larger than the ionization cross-section (see equation 5.12 in P. Hut & J. N. Bahcall (1983)), and therefore serves as an upper bound.
The probability that a binary formed at time $t_{\rm form}$ survives unionized and without experiencing a single encounter until the cluster evaporates at $t=t_{\rm evap}$ is given by
$P_{\rm no\,enc}=\exp\left(-\int_{t_{\rm form}}^{t_{\rm evap}}\Gamma_{\rm enc}(t)\mathrm{d}t\right),\ \Gamma_{\rm enc}=4\sqrt{\pi}n\sigma a^{2},$ (A14)
integrated along the binary’s orbit, where $\Gamma_{\rm enc}$ is the local encounter rate. The rate depends on the radius of the binary’s orbit in the cluster, $r(t)$, which itself varies with time due to the cluster’s evaporation. This process occurs on much longer time-scales than the orbit of a binary in a cluster (M. Gieles et al., 2011), so the orbit’s actions are conserved. One of these actions is the angular momentum of the orbit, which, for a circular orbit is given by $l=2m\sqrt{GM_{\rm cl}({<}r)\,r}$. Thus, for such an orbit
$r(t_{\rm form})M_{\rm cl}({<}r(t_{\rm form});t_{\rm form})=r(t)M_{\rm cl}({<}r(t);t),$ (A15)
and the orbital radius expands to compensate for the lost mass of the cluster. As the radial action (which measures the radial fluctuations of the orbit; D. Lynden-Bell 1963; J. Binney & S. Tremaine 2008) is also conserved, the eccentricity of the binary’s orbit about the cluster (defined as in D. Lynden-Bell 1963) cannot change. For highly eccentric orbits, as well as for nearly circular ones, the periapsis distance $r_{\min}$ is proportional to the angular momentum. Therefore, we may take Equation (A15) to imply that the pericenter of the orbit grows with time, because the enclosed mass $M_{\rm cl}(<r)$ decreases; a converse process also occurs in the adiabatic contraction of halos (J. Barnes & S. D. M. White, 1984; R. Jesseit et al., 2002).
Figure 13: The probability a binary will experience no strong encounter for the rest of a clusters’ life (Equation A16) versus the radial location in the cluster at the time the binary was born. The probability curves are organized such that columns correspond to initial cluster mass, $M_{\rm cl0}$, and rows correspond to the time since the cluster was born, $\tilde{\tau}$, in units of the total cluster lifetime. Binaries assembled with smaller SMA, later in a clusters life, or in lower mass clusters all have an elevated survival probability.
The encounter rate $\Gamma_{\rm enc}$ is proportional to the local number density, $n(r(t))$, and the local velocity dispersion, $\sigma(r(t))$, and thus depends on time via Equation (A15). The quantity $n\,\sigma$ is a decreasing function of $r$, so $\Gamma_{\rm enc}(r_{\min})$ is the maximum value of the encounter rate along an orbit. As the orbit expands, this upper bound on $\Gamma_{\rm enc}$ decreases. There are no explicit angle-action variables for a Plummer model, prohibiting the construction of an analytic formula for the time-dependent encounter rate $\Gamma_{\rm enc}(t)$. Instead, we seek a lower bound on the survival probability, $P_{\textrm{no enc}}$. Replacing $\Gamma_{\rm enc}(t)$ by its maximum initial value, $\Gamma_{\rm enc}(t_{\rm form})$, the survival probability becomes
$P_{\rm no\,enc}$ $\geq\exp\left[-\Gamma_{\rm enc}(t_{\rm form})\,(\tau_{\rm cl}-t_{\rm form})\right]$ (A16)
$\approx\exp\left[-4\sqrt{\pi}\,n(r_{\rm form},t_{\rm form})\,\sigma(r_{\rm form},t_{\rm form})\,a^{2}\,(\tau_{\rm cl}-t_{\rm form})\right].$
In Figure 13, $P_{\rm no\,enc}$ versus distance from cluster center, $r$, initial cluster masses ranging from $10^{2}$ to $10^{3.5}M_{\odot}$ and several key epochs in their evolution is displayed. For simplicity, we set the cluster orbital radius about the MW to be $R_{G}=8\,\rm{kpc}$, because most solar neighborhood 3BBs predicted by our model were emitted by clusters within $\pm 2\rm kpc$ of the Sun’s orbital radius. As expected, larger SMA have a smaller chance of surviving, while binaries assembled later in the clusters’ life are less likely to be perturbed before dissolution.
Encounters between two wide binaries are deemed negligible in our models because the binary-single encounter rate for wide binaries is always much larger than the wide-binary formation rate. This may be explicitly illustrated by, first, comparing the binary-binary and binary-single encounter rates,
$\frac{\Gamma_{\rm bb}}{\Gamma_{\rm enc}}\sim\frac{n_{\rm bin}\Sigma_{\rm bb}}{n\Sigma_{\rm enc}}\,.$ (A17)
The cross-section for a strong binary–binary encounter is approximately $\Sigma_{\rm bb}\sim\pi(a_{1}+a_{2})^{2}$ (J. M. Fregeau et al., 2004), while the cross-section for a strong binary-single encounter is $\Sigma_{\rm enc}\sim\pi a_{1}^{2}$. The steady-state number of wide binaries is suppressed relative to the number of single stars by a factor of the cluster size, i.e. $n_{\rm bin}/n\sim 1/N(t_{\rm form})$, where $N(t_{\rm form})$ is the number of bodies in the host cluster at binary formation (see J. Binney & S. Tremaine 2008, appendix M; M. Rozner & H. B. Perets 2023). Taking the SMA to be similar, we find
$\frac{\Gamma_{\rm bb}}{\Gamma_{\rm enc}}\sim\frac{4}{N(t_{\rm form})}\ll 1\,.$ (A18)
This relation approaches unity as the number of bodies in the cluster approaches $4$, at which point it is difficult to justify evolving the remaining system as a cluster or self-consistently calculating encounter rates.
Encounters between wide and hard binaries, of which there can be many, may be treated identically to that with a single star because a wide binary’s center of mass cannot significantly perturb a hard binary. Thus, we consider encounters between wide and hard binaries to be degenerate with binary-single encounters and included self-consistently within a Plummer cluster’s density profile.
@@ Appendix B Star Formation History and Radial Diffusion in the Milky Way’s Disks
Up until § A.5, we only considered dynamics internal to toy Milky Way stellar clusters. To estimate the number and orbital element distributions of 3BB’s in the solar neighborhood, it is critical to model the relationship between the clusters hosting 3BBF, the MW disk, the cluster birth time, $t$, and Galactocentric radius, $R_{G}$, in the Milky Way thin and thick disks, and the journey (i.e., angular momentum diffusion) the few 3BB’s managing to escape their home clusters must take to reach a comparatively small spatial region within ${\sim}100\,\rm{pc}$ of the Sun.
@@ B.1 Milky Way Disk Evolution
[…]
@@
where $R_{d0}=2.3\,\rm kpc$ is the disk scale-length, $R_{G}$ is the Galactocentric distance of the cluster’s orbit and the starting point from which escaping 3BBs begin their disk migration, $t$ is the time separation between the MW’s birth (assumed to occur $\tau_{\rm MW}=12\,\rm Gyrs$ ago) and the cluster’s birth, $\tau$ is the time between cluster birth and formation of the binary, $t^{\prime}=\tau_{\rm MW}-\tau_{\rm cl}-t$ is the look-back time since the host cluster of a 3BB dissolves, and $R_{G}^{\prime}$ is the location of the binary at $t^{\prime}$. Additionally, $\tau_{SFR}=6.8\,\rm Gyrs$ is the star formation timescale of the thick disk, $\tau_{ha}=6\,\rm Gyrs$ is the time when new cluster formation is “turned off” in the thick disk, $\tau_{m}=6\,\rm Gyrs$ is a timescale controlling the diffusion strength in the disk, and $c_{1}=2.48\,\rm{kpc}$ is the diffusion scale length. For simplicity, the latter two quantities and the form of the diffusion distribution, $D_{ha}$, are assumed identical to that of the thin disk models given by N. Frankel et al. (2020).
@@
[…]
@@
where $\tau_{la}=6\,\rm\,Gyr$ and $\tau_{SFR}=1\,\rm\,Gyr$. T. Wagg et al. (2022) constructed their Milky Way model assuming a total stellar mass $M_{d}=2.6\times 10^{10}M_{\odot}$ in each of the thick and thin disks. The number of stars in each present-day disk are $N_{\rm ha}=\frac{M_{d}}{m}=7\times 10^{10}$ and $N_{\rm la}=\frac{M_{d}}{m}=5.2\times 10^{10}$, respectively. By combining the two disk distributions, our evolving MW disk model is